find x:
log4(x+1)-log4(x-2)=1

Answer by edjones(7311)

log[4](x+1)-log[4](x-2)=1
log[4]((x+1)/(x-2))=1
4^log[4]((x+1)/(x-2))=4^1
(x+1)/(x-2)=4
x+1=4(x-2)
x+1=4x-8
-3x=-9
x=3
.
Ed